A regression line for predicting the selling prices of homes in Chicago is

ModifyingAbove y with caretyequals=168plus+​102x,

where x is the square footage of the house. A house with 1500 square feet recently sold for​ $140,000. What is the residual for this​ observation?
Answer #1

The regression equation: ŷ = 168 + 102x

The predicted price: ŷ = 168 + 102(1500) = 153168

The actual price: y = 140000

Thus, the residual for this observation

= Actual(y) - Predicted(ŷ)

= 140000 - 153168

= -13168